OpenStudy (anonymous):

Find the local extrema of f on [0, 2pi] and the subintervals on which f is increasing or is decreasing. f(x) = 2cosx + sin 2x

OpenStudy (anonymous):

f' = -2sinx+2cos2x = 0 =>sinx = cos2x => pi/2 - x = 2n(pi) + 2x =>-pi(4n-1)/6 = x

OpenStudy (anonymous):

note:the set ive given might be a subset of the actual solution set.since i've forgotten the cos equivalence conditions

myininaya (myininaya):

\[f'(x)=-2sinx+2\cos(2x)\] \[f'=0\] \[-2sinx+2\cos(2x)=0\] ----------------------------------------- Think: remember cos(2x)=cos(x+x)=cosx*cosx-sinx*sinx=cos^2x-sin^2x also remember sin^2x+cos^2x=1 so cos^2x=1-sin^2x so we can write cos(2x)=1-sin^2x-sin^2x=1-2sin^2x ----------------------------------------- so we have \[-2sinx+2(1-2\sin^2x)=0\] \[-2sinx+2-4\sin^2x=0\] \[-4\sin^2x-2sinx+2=0\] if you want it might be easy for you to write sin^2x as (sinx)^2 \[-4(sinx)^2-2sinx+2=0\] let u=sinx so u^2=sin^2x \[-4u^2-2u+2=0\] \[2u^2+u+1=0\] so we need to find u and then we can find x

myininaya (myininaya):

oops i made an error it is \[2u^2+u-1=0\]

myininaya (myininaya):

\[u=\frac{-1 \pm \sqrt{1-4(2)(-1)}}{2(2)}=\frac{-1 \pm \sqrt{1+8}}{4}=\frac{-1 \pm \sqrt{9}}{4}=\frac{-1 \pm 3}{4}\]

myininaya (myininaya):

\[u=\frac{-1+3}{4}=\frac{1}{2}, u=\frac{-1-3}{4}=-1\]

myininaya (myininaya):

remember u=sinx \[sinx=\frac{1}{2}, sinx=-1\]

myininaya (myininaya):

first equation gives \[x=\frac{\pi}{6}+2n \pi,\frac{5 \pi}{6}+2n \pi \] second equation gives \[x=\frac{3\pi}{2}+2n \pi\] for n=...,-2,-1,0,1,2,...

myininaya (myininaya):

but since we are only looking at 0 to 2pi we have x=pi/6,5pi/6, 3pi/2

myininaya (myininaya):

so there are numbers where we have horizontal tangents

myininaya (myininaya):

|--------|-------|------|---------| 0 pi/6 5pi/6 3pi/2 2pi plug in numbers between each critical number and endpoint (except before the first endpoint and after the last endpoint)

myininaya (myininaya):

any questions?

myininaya (myininaya):

f'(.5)=+ (we can actually just plug in 0 instead) f'(pi/2)=- f'(pi)=+ f'(11pi/6)=+ so increasing on intervals: (0,pi/6)U(5pi/6,3pi/2)

myininaya (myininaya):

so at pi/6 we had it went from increasing before to decreasing after so we have a max occuring at pi/6 we have a min occuring at 5pi/6 since it goes from decreasing to increasing

myininaya (myininaya):

oh yeah and it is decreasing on the interval (pi/6,5pi/6)